OpenStudy (agent_a):

Calculus Question: I'm in the middle of solving a problem, although I was wondering how: (hold on, I'm typing it in the equation formatter)

OpenStudy (agent_a):

\[-\sin^2(t)+\cos^2(t)\] simplifies to: \[\cos(2t)\]

OpenStudy (agent_a):

Similarly, \[-4\sin^2(t)+4\cos^2(t)\] Simplifies to: \[4\cos(2t)\] How did that happen?

OpenStudy (agent_a):

(Done typing question)

OpenStudy (cwrw238):

this is a known trig identity cos2t = cos^t - sin^2t

OpenStudy (agent_a):

Oh of course.... I didn't think of that. Thanks a lot!